OpenStudy (anonymous):

In a white water relay race, one kayak team averaged 26km/h downstream but only 6km/h upstream in the white water course. What was the rate of the kayak in still water? What was the rate of the current?(solve using linear systems )

OpenStudy (anonymous):

let v be the speed of the kayak and x be the speed of the current v+x=26 v-x=6 2x=20 x=10 v=16

OpenStudy (anonymous):

Thanks its working expect where did the 20 come from?? @carl_gauss

OpenStudy (anonymous):

I. v+x=26 II. v-x=6 Gaussian elimination: I-II

OpenStudy (anonymous):

carl friedrich gauss

OpenStudy (anonymous):

indeed xD

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Ok then... But then wouldn't the other variable be 14 ... 6-20... -14

OpenStudy (anonymous):

@carl_gauss ??

OpenStudy (anonymous):

no the speed of the boat is 16 and the current is 10 if the boat drives with the current it has a speed of 26 (16+10) if the boat drives against the current it's speed is 6 (16-10) if there is no current it has just a speed of 16. i'll write more: I. v+x=26 II. v-x=6 I-II=III: (v-v)+(x-(-x))=26-6 III: 0+2x=20 |/2 III: x=10 the twenty just is a interim result and not used for the speeds itself

OpenStudy (anonymous):

Okay thanks !!!